OpenStudy (anonymous):

A wombat runs south in a straight line with an average velocity of 5 m/s for 4 minutes and then with an average velocity of 4 m/s for 3 minutes in the same direction . Would those both be the initial and Final velocities ?

OpenStudy (sasogeek):

does the question states if it comes to a stop? if it doesn't state anything of the sort then I'd say yes and I'll solve for whatever with those as initial and final velocities

OpenStudy (anonymous):

Btw what is the displacement is what the question is

OpenStudy (anonymous):

asking

OpenStudy (anonymous):

Yeah, it doesn't mention anything about stopping

OpenStudy (sasogeek):

displacement is velocity times time taken

OpenStudy (anonymous):

I got -3.5?

OpenStudy (sasogeek):

convert minutes to seconds then work it out

OpenStudy (sasogeek):

i got 1920m as distance due south = displacement

OpenStudy (anonymous):

Ohhh I missed that . Now I just got -210

OpenStudy (sasogeek):

how is displacement negative? how did u work it out?

OpenStudy (sasogeek):

there's no such thing as negative displacement or distance

OpenStudy (anonymous):

I did final veloctiy - initial velocity * time / 2

OpenStudy (sasogeek):

why?

OpenStudy (anonymous):

Our teacher told us to use this long equation. Just tell me how you would work it?

OpenStudy (sasogeek):

ok we're looking for displacement... right? in that case, what is displacement... speed x time. Initially, the average speed was 5m/s and that lasted a period of 4 minutes then average speed after 4 minutes is 4m/s for a period of 3 minutes. We have speed, we have time... But our units don't match, hence... 4minutes=240seconds 3minutes=180seconds Now we can safely find total displacement = (5*240) + (4*180) =1200+720 =1920 I hope that made some sense :/

OpenStudy (sasogeek):

sorry i forgot the units.... 1920m due south

OpenStudy (anonymous):

Yes that did make sense ! thank you! I have one more thing to ask

OpenStudy (sasogeek):

k

OpenStudy (anonymous):

This problem says a falling skydiver slows from a speed of 52 m/s to 8m/s in .8 sec as the parachute opens. Would if make sense to get a change in velocity that is negative since he is slowing down? Are those even velocities?

OpenStudy (sasogeek):

hmm... the change in velocity won't be negative, I'm saying this cos I can't figure out why it should be negative but it could be... i won't conclude on this but I'll let u know once i find out what the right thing is

OpenStudy (anonymous):

Alright alright

OpenStudy (sasogeek):

what is change in velocity?

OpenStudy (sasogeek):

final velocity - initial velocity.... now tell me what u think?

OpenStudy (sasogeek):

i think i have my conclusion... and it is yes, you could say there was a negative velocity change there

OpenStudy (anonymous):

I thought so! Thanks man you were a great deal of help! Hopefully I can ace my test tomorrow!

OpenStudy (sasogeek):

I could say the change in velocity was 44m/s... but was the velocity increasing or decreasing? the negative gives us an idea of whether the velocity was increasing or decreasing so yeah ;) that's my conclusion :D

OpenStudy (anonymous):

Alright sounds good! :D Thank you so much!!!

OpenStudy (sasogeek):

anytime :) you're welcome :D check out my website http://sasogeek.com i cleaned it off everything and i'm reconstructing it, when it is ready and I put the chat room there, i'll let you know to pass by sometime to say hi :)

OpenStudy (anonymous):

Awesome! I'll check it forsure!

OpenStudy (sasogeek):

hey I wanna make a youtube channel and put math and physics videos there... wanna help out?

OpenStudy (sasogeek):

as well as tech videos ;)